An ideal gas turbine operates using air (k=1.4) coming at 375°C and 300 kPa. Find the exit pressure (kPa)  if the air exits at 125 oC. Assume the air behaves ideally, properties are constant, and the turbine operates isentropically.

The correct answer is 30.8 kPa
Answer #1

\\ \textit{\textbf{Solution:}}\\ \\ \qquad \textup{Given Data :-}\\ \\ \qquad T_1=375^oC=648\, \, K \\ P_1=300\, \, kPa \\ T_2 = 125^oC =398\, \, K \\ k=1.4\\ \\ \textup{For an ideal gas under isentropic relation and constant specific } \\ \textup{heat assumption, the temperature ratio is written as:}\\ \\ \therefore \frac{T_2}{T_1}=\left (\frac{P_2}{P_1} \right )^{(k-1)/k} \\ \\ \therefore P_2=P_1 \left ( \frac{T_2}{T_1} \right )^{\frac{k}{k-1}} \\ \\ \therefore P_2=300\left(\frac{398}{648}\right)^{\frac{1.4}{1.4-1}} \\ \\ {\color{DarkBlue} \boxed {\therefore P_2= 54.4752\, \, kPa} }\\

\textup{(Note : The correct answer is 54.4752 kPa and the answer}\textup{mentioned in problem statement is incorrect.)}